1. If one glucose molecule enters into glycolysis and proceeds through the PDC to acetyl-CoA, how many NET ATP are generated in these reactions?

("Net" refers to total # generated - total number consumed.)

2. If one glucose molecule enters into glycolysis and proceeds through the PDC to acetyl-CoA, how many CO2 molecules are released in these reactions?

3. If one glucose molecule enters into glycolysis and proceeds through the PDC to acetyl-CoA, how many NADH molecules are generated in these reactions?
Answer #1

During glycolysis, one molecule of glucose gives 2 molecules of Pyruvate. This Pyruvate proceeds through the PDC to acetyl-CoA.

1. 2ATP used during glycolysis, 4 are produced by substrate level phosphorylation. 6 ATP are produced by oxidative phosphorylation.

Through PCD, 6 ATP produced by oxidative phosphorylation (considering 2 molecules of pyruvate).

So total generated is 16 ATP and consumed is 2 ATP.

SO, Net ATP = 16-2= 14 ATP

2. There is no CO2 evolution during glycolysis but one pyruvate converts into Acetyl co-A by giving one CO2 molecule. So total 2 CO2 molecules are released.

3. 2 NADH molecules are produced during glycolysis. Each pyruvate gives one NADH. So net NADH generation is 4 NADH.

Glucose ATP ADP a G-6-P F-6-P ATP - ADP & F-1,6-bip 2X PGAL 2x 1,3-Di PGAL 2 ZNAD ZNADH = GATP (by osidative phosphorylation

2x Pyruvate 2 NAD POC - ZNADH₂ = G A TP (Oreo. Phosp.) 2 coz 2x Acetyl Co-A